OpenStudy (anonymous):

three cube roots of 8(1-j) in polar form

OpenStudy (kc_kennylau):

Rewrite it in r(cosθ+jsinθ) first

OpenStudy (kc_kennylau):

The roots would be r^(1/3)(cos(θ/3)+jsin(θ/3)), r^(1/3)(cos((θ+2pi)/3)+jsin((θ+2pi)/3)), and r^(1/3)(cos((θ+4pi)/3)+jsin((θ+4pi)/3))

OpenStudy (kc_kennylau):

\[\Large\sqrt[3]r\left(\cos\frac\theta3+j\sin\frac\theta3\right)\]\[\Large\sqrt[3]r\left(\cos\frac{\theta+2\pi}3+j\sin\frac{\theta+2\pi}3\right)\]\[\Large\sqrt[3]r\left(\cos\frac{\theta+4\pi}3+j\sin\frac{\theta+4\pi}3\right)\]